OpenStudy (anonymous):

***PLEASE HELP***SUPER EASY*** 1/189 as a decimal?

OpenStudy (anonymous):

I Will give a medal

OpenStudy (anonymous):

can you help me here? Pint-Sized Portraits has 5 different backgrounds and 3 poses in which to photograph children. How many different pictures are there to choose from? Ans: 8,15 or 24

OpenStudy (anonymous):

if its super easy... why not do it yourself??? 0.o

OpenStudy (anonymous):

@Guitarestfreak to find however many pictures/combinations to choose from just multiply. So, for this question it would be 5*3 =?. 5*3 = 15 so the answer would be 15

OpenStudy (anonymous):

thanks!!!

OpenStudy (anonymous):

No problem!

OpenStudy (anonymous):

Evaluate 3! + 4!.

OpenStudy (anonymous):

30, 144, or 5,040?

OpenStudy (anonymous):

I don't know how to explain it totally but is it ok if I just give you the answer without showing the steps?

OpenStudy (anonymous):

@Guitarestfreak

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

Ok so 3! = 6, and 4! = 24. So 3! + 4! would be like 6+24 = ?. 6+24 = 30

OpenStudy (anonymous):

Any other questions? @Guitarestfreak

OpenStudy (anonymous):

lol lots

OpenStudy (anonymous):

Ok I'll try to answer some... @Guitarestfreak

OpenStudy (anonymous):

let me write something down

OpenStudy (anonymous):

@Guitarestfreak

OpenStudy (anonymous):

next one

OpenStudy (anonymous):

@Guitarestfreak Ok

OpenStudy (anonymous):

if a salesperson wanted to give the impression that there was a larger increase in sales over the time period. what he might do to the graph? Sry I can't do the graph on here:( Ans: use a smaller rang, use a larger range, or delete a month??

OpenStudy (anonymous):

What type of graph is it? Is it a histogram, double bar graph, etc?

OpenStudy (anonymous):

@Guitarestfreak

OpenStudy (anonymous):

back

OpenStudy (anonymous):

i think histogram

OpenStudy (anonymous):

ill just skip that one

OpenStudy (anonymous):

in how many different ways can 5 books be arranged on a shelf? Ans: 5, 25, or 120?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Ok so the answer would be 120

OpenStudy (anonymous):

Because since you have 5 books you would do it like this: 5*4*3*2*1 = 120. Get it? @Guitarestfreak

OpenStudy (anonymous):

another one too. The mean weight of the packages Joan shipped was 2.5 pounds. If John mailed four packages and three of them had weights of 1.7, 3.2 and 2.8 pounds, then what did the other package weigh? Ans: 2.3, 2.8, or 3.3?

OpenStudy (anonymous):

so x those numbers?

OpenStudy (anonymous):

Let me see.

OpenStudy (anonymous):

i think you're right there, thanks :)

OpenStudy (anonymous):

The mean weight of the packages Joan shipped was 2.5 pounds. If John mailed four packages and three of them had weights of 1.7, 3.2 and 2.8 pounds, then what did the other package weigh? Ans: 2.3 ANSWER <-----

OpenStudy (anonymous):

thanks!! sounds right.

OpenStudy (anonymous):

how did u figure it out?

OpenStudy (anonymous):

Evaluate C(n,n). ans: 0, 1, or n?

OpenStudy (anonymous):

Wel, I decided to try all the numbers to see which one included in the group of numbers has 2.5 as the mean. To calculate mean you add up all the numbers and divide it by however many numbers there are.

OpenStudy (anonymous):

I don't know the answer to that question above sorry! @Guitarestfreak

OpenStudy (anonymous):

kk that helps.

OpenStudy (anonymous):

what about this one? you want to know the brand of jeans worn most often by people 20 to 30 years old. which of the following surveys is biased? Ans: surveying people 20 to 30 years old at a store that sells one brand of jeans, surveying people 20 to 30 old at a store that sells only jeans, or surveying everyone at a store that sells only jeans??

OpenStudy (anonymous):

surveying people 20 to 30 years old at a store that sells one brand of jeans, @Guitarestfreak

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

and this one too

OpenStudy (anonymous):

the chorus has 30 members. Which expression represents the number of ways a group of 6 members can be chosen to do a special performance? Ans: 

OpenStudy (anonymous):

sry its crazyy lol

OpenStudy (anonymous):

30/6!24!

OpenStudy (anonymous):

another one. Which of the following values is the 75th percentile? Ans: lower quartile, median, upper quartile?

OpenStudy (anonymous):

Upper quartile I think

OpenStudy (anonymous):

k, brb

OpenStudy (anonymous):

how many different arrangements of the letters in the word "COLOR" are there? Ans: 20, 60, 120? brb, but u can continue to anwser this question

OpenStudy (anonymous):

@Guitarestfreak 120

OpenStudy (anonymous):

@mangorox, it isn't a nice decimal. why do you need it as a decimal?